OpenStudy (anonymous):

write the equation of a line given the points (1,1) (4,16)

OpenStudy (anonymous):

do you know how to find slope?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(1\quad ,&1)\quad &(4\quad ,&16) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}\)